OpenStudy (anonymous):

(n+4)!/(n²+7n+12)n! solve please

OpenStudy (anonymous):

n/n+3

OpenStudy (loser66):

@Jpenrod how to get it?

OpenStudy (anonymous):

I just simplified Idk if its right..

OpenStudy (cwrw238):

n^2 + 7n + 12 = (n+4)(n+ 3) (n+ 4) ! / n! = (n + 4)(n + 3)(n + 2)(n + 1) can you continue?

OpenStudy (loser66):

Thanks @cwrw238

OpenStudy (cwrw238):

ok